# Stability of Poiseuille Flow of Navier-Stokes Equations on $\mathbb{R}^{2}$

### v1

@@ 3 Non-linear Problem
For the non-linear problem, the non-linear term $(\nabla^{\bot}\psi\cdot\nabla\omega)_{k}$ combines multiple frequencies. Hence the energy $E_{k}$ that depends on $k$ is not sufficient. Thus we define the following energies and dissipations.
@@
$\mathcal{E}=\mathcal{E}_{1}+\mathcal{E}_{2}=\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}E_{k}\mathrm{d}k+\sup_{k}\frac{1}{2}\|y\omega_{k}\|_{2}^{2}+\|\nabla_{k}\psi_{k}\|_{2}^{2},$
@@
$\mathcal{D}=\mathcal{D}_{1}+\mathcal{D}_{2}=\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}D_{k}\mathrm{d}k+\sup_{k}\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}+\nu\|\omega_{k}\|_{2}^{2},$
where $J\in[1,+\infty),m\in(\frac{3}{4},1)$, $M_{k}(t)$ solves the ODE
@@
The multiplier $M_{k}(t)$ is included in $\mathcal{E}$ to address terms in $\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}$ which arise from the time-derivative falling on $\langle c\lambda_{k}t\rangle^{2J}$. Solving the ODE explicitly, we can get $1\leq M_{k}(t)\leq e^{\frac{4J^{2}}{3}}$ for all $t\geq 0$ and all $k\in\mathbb{R}$. Thus $M_{k}(t)$ is uniformly bounded above and below in both $k$ and $t$.
@@
$NL_{k}=-(\nabla^{\bot}\psi\cdot\nabla\omega)_{k}=-\int\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\cdot\nabla_{k^{\prime}}\omega_{k^{\prime}}\mathrm{d}k^{\prime},$
then $\partial_{t}\omega_{k}=L_{k}+NL_{k}$. We divide $\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}$ into the linear part, the nonlinear part, and the terms where the time derivative acts on the multipliers.
$\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}=\mathcal{L}+\mathcal{NL}+\int Jc\lambda_{k}\frac{2c\lambda_{k}t}{\langle c\lambda_{k}t\rangle^{2}}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}E_{k}\mathrm{d}k-\int\frac{M_{k}^{\prime}(t)}{M_{k}(t)}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}E_{k}\mathrm{d}k,$
@@
$\mathcal{L}=$ $\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\bigg{(}\mathrm{Re}\langle\omega_{k},L_{k}\rangle+c_{\alpha}\alpha_{k}\mathrm{Re}\langle\nabla_{k}\omega_{k},\nabla_{k}L_{k}\rangle+2c_{\beta}\beta_{k}(\mathrm{Re}\langle ikyL_{k},\partial_{y}\omega_{k}\rangle\mathrm{Re}\langle iky\omega_{k},\partial_{y}L_{k}\rangle)$
$+c_{\gamma}\gamma_{k}(\mathrm{Re}\langle y\omega_{k},yL_{k}\rangle+2\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}L_{k}\rangle)\bigg{)}\mathrm{d}k+\sup_{k}\mathrm{Re}\langle y\omega_{k},yL_{k}\rangle+2\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}L_{k}\rangle,$
@@
$\mathcal{NL}=$ $\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\bigg{(}\mathrm{Re}\langle\omega_{k},NL_{k}\rangle+c_{\alpha}\alpha_{k}\mathrm{Re}\langle\nabla_{k}\omega_{k},\nabla_{k}NL_{k}\rangle+2c_{\beta}\beta_{k}(\mathrm{Re}\langle ikyNL_{k},\partial_{y}\omega_{k}\rangle+\mathrm{Re}\langle iky\omega_{k},\partial_{y}NL_{k}\rangle)$
@@
$+c_{\gamma}\gamma_{k}(\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle+2\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle)\bigg{)}\mathrm{d}k+\sup_{k}\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle+2\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle.$
By Lemma 2.3 and Theorem 2.1, we have
@@
$\mathcal{L}\leq\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}(-4cD_{k}-4c\lambda_{k}E_{k})\mathrm{d}k+\sup_{k}-\mathcal{D}_{2}\leq-4c\mathcal{D}-4\int c\lambda_{k}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}E_{k}\mathrm{d}k,$
since $c$ is sufficiently small. By Young’s inequality and the definition of $M_{k}(t)$,
$\int Jc\lambda_{k}\frac{2c\lambda_{k}t}{\langle c\lambda_{k}t\rangle^{2}}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}E_{k}\mathrm{d}k\leq\int c\lambda_{k}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}E_{k}\mathrm{d}k+\int\frac{M_{k}^{\prime}(t)}{M_{k}(t)}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}E_{k}\mathrm{d}k.$
We observe that
$\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}\leq-4c\mathcal{D}+\mathcal{NL}.$ (3.1)
Thus it suffices to give the bound of $\mathcal{NL}$. In order to estimate $\mathcal{NL}$, we separate it into the following terms
$T_{1}=\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\mathrm{Re}\langle\omega_{k},NL_{k}\rangle\mathrm{d}k,$
$T_{2}=c_{\alpha}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\alpha_{k}\mathrm{Re}\langle\nabla_{k}\omega_{k},\nabla_{k}NL_{k}\rangle\mathrm{d}k,$
$T_{3}=2c_{\beta}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle ikyNL_{k},\partial_{y}\omega_{k}\rangle\mathrm{d}k,$
$T_{4}=2c_{\beta}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle iky\omega_{k},\partial_{y}NL_{k}\rangle\mathrm{d}k,$
$T_{5}=c_{\gamma}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle\mathrm{d}k,$
$T_{6}=2c_{\gamma}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle\mathrm{d}k,$
$T_{7}=\sup_{k}\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle,$
@@
$T_{8}=2\sup_{k}\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle.$
@@ 3.1 Technical Lemmas
Now, we give some minor technical lemmas will be used throughout the eatimates of $\mathcal{NL}$. The proofs rely on basic inequalities of analysis, namely Hölder’s inequality and Gagliardo-Nirenberg-Sobolev inequality.
@@
$\int\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k\lesssim\nu^{-\frac{5}{6}}\mathcal{D}^{\frac{1}{2}},$ (3.2)
$\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\mathcal{E}^{\frac{1}{2}},$ (3.3)
$\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}.$ (3.4)
@@
$\int\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k\lesssim\int\|\nabla_{k}\omega_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\nabla_{k}\omega_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k\\ \lesssim\int\|\nabla_{k}\omega_{k}\|_{2}^{\frac{1}{2}}\|\Delta_{k}\omega_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k.$
@@
$\int\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k\lesssim\left(\int\langle k\rangle^{-2m}\alpha_{k}^{-\frac{1}{2}}\nu^{-1}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle k\rangle^{2m}\nu\|\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{4}}\left(\int\langle k\rangle^{2m}\alpha_{k}\nu\|\Delta_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{4}}.$
For the first term, we split the integral domain into $|k|\geq\nu^{-\frac{1}{3}}$ and $|k|\leq\nu^{-\frac{1}{3}}$, then the $|k|\geq\nu^{-\frac{1}{3}}$ term obeys
@@
where we have used $m>\frac{3}{4}$ and $\nu<1$. Then we integrate over $|k|\leq\nu^{-\frac{1}{3}}$,
$\int_{|k|\leq\nu^{-\frac{1}{3}}}\langle k\rangle^{-2m}\alpha_{k}^{-\frac{1}{2}}\nu^{-1}\mathrm{d}k\lesssim\int_{|k|\leq\nu^{-\frac{1}{3}}}\left(\nu^{\frac{2}{3}}\right)^{-\frac{1}{2}}\nu^{-1}\mathrm{d}k\approx\nu^{-\frac{5}{3}},$
Combining the two part and recalling the definition of $\mathcal{D}$, then (3.2) follows. The proof of (3.3) is similar to (3.2). Indeed
$\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k$ $\lesssim\int\|\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k\lesssim\int\|\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\|\omega_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k$
$\lesssim\left(\int\langle k\rangle^{-2m}\gamma_{k}^{-\frac{1}{2}}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle k\rangle^{2m}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{4}}\left(\int\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{4}}.$
We still split the integral domain into $|k|\geq\nu^{-\frac{1}{3}}$ and $|k|\leq\nu^{-\frac{1}{3}}$, and recall $m>\frac{3}{4}$, then we have
@@
$\int_{|k|\leq\nu^{-\frac{1}{3}}}\langle k\rangle^{-2m}\gamma_{k}^{-\frac{1}{2}}\mathrm{d}k\lesssim\int_{|k|\leq\nu^{-\frac{1}{3}}}\left(\nu^{-\frac{2}{3}}\right)^{-\frac{1}{2}}\mathrm{d}k\approx 1.$
Hence (3.3) follows. Turning to (3.4), we use Gagliardo-Nirenberg-Sobolev inequality,
$\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\int|k|^{-\frac{1}{2}}|k|^{\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k\lesssim\int|k|^{-\frac{1}{2}}\|\Delta_{k}\psi_{k}\|_{2}\mathrm{d}k=\int|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k.$
For the integral over $|k|\geq\nu^{-\frac{1}{3}}$, by Hölder’s inequality, we have
$\int_{|k|\geq\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle k\rangle^{-2m}(\gamma_{k}\nu)^{-1}|k|^{-1}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle k\rangle^{2m}\gamma_{k}\nu\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\lesssim\mathcal{D}^{\frac{1}{2}},$
so (3.4) follows from
$\int_{|k|\leq\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|\omega_{k}\|_{2}\mathrm{d}k\leq\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\mathrm{d}k\sup_{k}\|\omega_{k}\|_{2}\lesssim\nu^{-\frac{1}{6}}(\nu^{-1}\mathcal{D})^{\frac{1}{2}}=\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}.\qed$

### v4

@@ 3 Non-linear Problem
For the non-linear problem, the non-linear term $-(\nabla^{\bot}\psi\cdot\nabla\omega)_{k}$ combines multiple frequencies. Hence the energy $E_{k}$ that depends on $k$ is not sufficient. Thus we define the following energies and dissipations.
@@
$\mathcal{E}=\mathcal{E}_{1}+\mathcal{E}_{2}=\int_{\mathbb{R}}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}E_{k}\mathrm{d}k+\sup_{k\in\mathbb{R}}\left(\frac{1}{2}\|y\omega_{k}\|_{2}^{2}+\|\nabla_{k}\psi_{k}\|_{2}^{2}\right),$
@@
$\mathcal{D}=\mathcal{D}_{1}+\mathcal{D}_{2}=\int_{0}^{t}\widetilde{\mathcal{D}}\mathrm{d}s+\sup_{k\in\mathbb{R}}\int_{0}^{t}\nu\|\omega_{k}\|_{2}^{2}+\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}s,$
where
$\widetilde{\mathcal{D}}=\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}D_{k}\mathrm{d}k,$
$J\in[1,+\infty),m\in(\frac{3}{4},1)$, $M_{k}(t)$ solves the ODE
@@
The multiplier $M_{k}(t)$ is included in $\mathcal{E}_{1}$ to address terms in $\frac{\mathrm{d}\mathcal{E}_{1}}{\mathrm{d}t}$ which arise from the time-derivative falling on $\langle c\lambda_{k}t\rangle^{2J}$. Solving the ODE explicitly, we can get $1\leq M_{k}(t)\leq e^{\frac{\pi J^{2}}{4}}$ for all $t\geq 0$ and all $k\in\mathbb{R}$. Thus $M_{k}(t)$ is uniformly bounded above and below in both $k$ and $t$.
@@
$NL_{k}=-(\nabla^{\bot}\psi\cdot\nabla\omega)_{k}=-\int_{\mathbb{R}}\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\cdot\nabla_{k^{\prime}}\omega_{k^{\prime}}\mathrm{d}k^{\prime},$
then
$\partial_{t}\omega_{k}=L_{k}+NL_{k}.$
By Newton-Leibniz formula,
$\mathcal{E}(t)+\mathcal{D}_{2}(t)=$ $\mathcal{E}_{1}(0)+\int_{0}^{t}\frac{\mathrm{d}\mathcal{E}_{1}}{\mathrm{d}s}\mathrm{d}s+\sup_{k\in\mathbb{R}}\left(\frac{1}{2}\|y\omega_{k}(0)\|_{2}^{2}+\|\nabla_{k}\psi_{k}(0)\|_{2}^{2}+\int_{0}^{t}\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{1}{2}\|y\omega_{k}\|_{2}^{2}+\|\nabla_{k}\psi_{k}\|_{2}^{2}\right)\mathrm{d}s\right)$
$+\sup_{k\in\mathbb{R}}\int_{0}^{t}\nu\|\omega_{k}\|_{2}^{2}+\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}s$
$\leq$ $\mathcal{E}_{1}(0)+\int_{0}^{t}\frac{\mathrm{d}\mathcal{E}_{1}}{\mathrm{d}s}\mathrm{d}s+2\sup_{k\in\mathbb{R}}\bigg{(}\frac{1}{2}\|y\omega_{k}(0)\|_{2}^{2}+\|\nabla_{k}\psi_{k}(0)\|_{2}^{2}+\int_{0}^{t}\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{1}{2}\|y\omega_{k}\|_{2}^{2}+\|\nabla_{k}\psi_{k}\|_{2}^{2}\right)\mathrm{d}s$
$+\int_{0}^{t}\nu\|\omega_{k}\|_{2}^{2}+\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}s\bigg{)}$
$\leq$ $2\mathcal{E}(0)+\int_{0}^{t}\frac{\mathrm{d}\mathcal{E}_{1}}{\mathrm{d}s}\mathrm{d}s+2\sup_{k\in\mathbb{R}}\int_{0}^{t}\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{1}{2}\|y\omega_{k}\|_{2}^{2}+\|\nabla_{k}\psi_{k}\|_{2}^{2}\right)+\nu\|\omega_{k}\|_{2}^{2}+\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}s.$ (3.1)
We divide $\frac{\mathrm{d}\mathcal{E}_{1}}{\mathrm{d}s}$ into the linear part, the nonlinear part, and the terms where the time derivative acts on the multipliers.
$\frac{\mathrm{d}\mathcal{E}_{1}}{\mathrm{d}s}=\mathcal{L}_{1}+\mathcal{NL}_{1}+\int_{\mathbb{R}}Jc\lambda_{k}\frac{2c\lambda_{k}s}{\langle c\lambda_{k}s\rangle^{2}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}E_{k}\mathrm{d}k-\int_{\mathbb{R}}\frac{M_{k}^{\prime}(s)}{M_{k}(s)}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}E_{k}\mathrm{d}k,$
@@
$\mathcal{L}_{1}=$ $\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\bigg{(}\mathrm{Re}\langle\omega_{k},L_{k}\rangle+c_{\alpha}\alpha_{k}\mathrm{Re}\langle\nabla_{k}\omega_{k},\nabla_{k}L_{k}\rangle+2c_{\beta}\beta_{k}(\mathrm{Re}\langle ikyL_{k},\partial_{y}\omega_{k}\rangle+\mathrm{Re}\langle iky\omega_{k},\partial_{y}L_{k}\rangle)$
$+c_{\gamma}\gamma_{k}(\mathrm{Re}\langle y\omega_{k},yL_{k}\rangle+2\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}L_{k}\rangle)\bigg{)}\mathrm{d}k,$
@@
$\mathcal{NL}_{1}=$ $\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\bigg{(}\mathrm{Re}\langle\omega_{k},NL_{k}\rangle+c_{\alpha}\alpha_{k}\mathrm{Re}\langle\nabla_{k}\omega_{k},\nabla_{k}NL_{k}\rangle+2c_{\beta}\beta_{k}(\mathrm{Re}\langle ikyNL_{k},\partial_{y}\omega_{k}\rangle+\mathrm{Re}\langle iky\omega_{k},\partial_{y}NL_{k}\rangle)$
@@
$+c_{\gamma}\gamma_{k}(\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle+2\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle)\bigg{)}\mathrm{d}k.$
By and Theorem 2.1, we have
@@
$\mathcal{L}_{1}\leq\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}(-4cD_{k}-4c\lambda_{k}E_{k})\mathrm{d}k=-4c\widetilde{\mathcal{D}}-4\int_{\mathbb{R}}c\lambda_{k}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}E_{k}\mathrm{d}k.$
Using Young’s inequality and recalling the definition of $M_{k}(t)$, we have
$\int_{\mathbb{R}}Jc\lambda_{k}\frac{2c\lambda_{k}s}{\langle c\lambda_{k}s\rangle^{2}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}E_{k}\mathrm{d}k$ $\leq\int_{\mathbb{R}}c\lambda_{k}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}E_{k}\mathrm{d}k+\int_{\mathbb{R}}cJ^{2}\lambda_{k}\frac{(c\lambda_{k}s)^{2}}{\langle c\lambda_{k}s\rangle^{4}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}E_{k}\mathrm{d}k$
$=\int_{\mathbb{R}}c\lambda_{k}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}E_{k}\mathrm{d}k+\int_{\mathbb{R}}\frac{M_{k}^{\prime}(s)}{M_{k}(s)}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}E_{k}\mathrm{d}k.$
Thus we observe that
$\frac{\mathrm{d}\mathcal{E}_{1}}{\mathrm{d}t}\leq-4c\widetilde{\mathcal{D}}+\mathcal{NL}_{1}.$
Inserting the above equation and (2.8), (2.9) from Lemma 2.3 into (3), we get
@@
$\mathcal{E}(t)+\mathcal{D}_{2}(t)\leq 2\mathcal{E}(0)-4c\mathcal{D}_{1}(t)+\int_{0}^{t}\mathcal{NL}_{1}\mathrm{d}s+2\sup_{k\in\mathbb{R}}\left(\int_{0}^{t}\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle+2\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle\mathrm{d}s\right),$
which implies
$\mathcal{E}(t)\leq 2\mathcal{E}(0)-4c\mathcal{D}(t)+\mathbb{NL}(t),$ (3.2)
since $c$ is sufficiently small, where
$\mathbb{NL}(t)=\int_{0}^{t}\mathcal{NL}_{1}\mathrm{d}s+2\sup_{k\in\mathbb{R}}\int_{0}^{t}\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle\mathrm{d}s+4\sup_{k\in\mathbb{R}}\int_{0}^{t}\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle\mathrm{d}s.$
Thus it suffices to give the bound of $\mathbb{NL}$. In order to estimate $\mathbb{NL}$, we separate it into the following terms
$T_{1}=\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\mathrm{Re}\langle\omega_{k},NL_{k}\rangle\mathrm{d}k\mathrm{d}s,$
$T_{2}=c_{\alpha}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\alpha_{k}\mathrm{Re}\langle\nabla_{k}\omega_{k},\nabla_{k}NL_{k}\rangle\mathrm{d}k\mathrm{d}s,$
$T_{3}=2c_{\beta}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle ikyNL_{k},\partial_{y}\omega_{k}\rangle\mathrm{d}k\mathrm{d}s,$
$T_{4}=2c_{\beta}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\beta_{k}\mathrm{Re}\langle iky\omega_{k},\partial_{y}NL_{k}\rangle\mathrm{d}k\mathrm{d}s,$
$T_{5}=c_{\gamma}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle\mathrm{d}k\mathrm{d}s,$
$T_{6}=2c_{\gamma}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle\mathrm{d}k\mathrm{d}s,$
$T_{7}=2\sup_{k\in\mathbb{R}}\int_{0}^{t}\mathrm{Re}\langle y\omega_{k},yNL_{k}\rangle\mathrm{d}s,$
$T_{8}=4\sup_{k\in\mathbb{R}}\int_{0}^{t}\mathrm{Re}\langle\nabla_{k}\psi_{k},\nabla_{k}\Delta_{k}^{-1}NL_{k}\rangle\mathrm{d}s.$
@@ 3.1 Technical Lemmas
Now, we give some minor technical lemmas will be used throughout the eatimates of $\mathbb{NL}$. The proofs rely on basic inequalities of analysis, namely Hölder’s inequality and Gagliardo-Nirenberg-Sobolev inequality.
@@
$\int_{\mathbb{R}}\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k\lesssim\nu^{-\frac{5}{6}}\widetilde{\mathcal{D}}^{\frac{1}{2}},$ (3.3)
$\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\mathcal{E}^{\frac{1}{2}},$ (3.4)
@@
$\int_{\mathbb{R}}\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k\lesssim\int_{\mathbb{R}}\|\nabla_{k}\omega_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\nabla_{k}\omega_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k\\ \lesssim\int_{\mathbb{R}}\|\nabla_{k}\omega_{k}\|_{2}^{\frac{1}{2}}\|\Delta_{k}\omega_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k.$
@@
$\int_{\mathbb{R}}\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k\lesssim\left(\int_{\mathbb{R}}\langle k\rangle^{-2m}\alpha_{k}^{-\frac{1}{2}}\nu^{-1}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle k\rangle^{2m}\nu\|\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{4}}\left(\int_{\mathbb{R}}\langle k\rangle^{2m}\alpha_{k}\nu\|\Delta_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{4}}.$
For the first term, we split the integral domain into $|k|\geq\nu^{-\frac{1}{3}}$ and $|k|<\nu^{-\frac{1}{3}}$, then the $|k|\geq\nu^{-\frac{1}{3}}$ term obeys
@@
where we have used $m>\frac{3}{4}$ and $\nu<1$. Then we integrate over $|k|<\nu^{-\frac{1}{3}}$,
$\int_{|k|<\nu^{-\frac{1}{3}}}\langle k\rangle^{-2m}\alpha_{k}^{-\frac{1}{2}}\nu^{-1}\mathrm{d}k\lesssim\int_{|k|<\nu^{-\frac{1}{3}}}\left(\nu^{\frac{2}{3}}\right)^{-\frac{1}{2}}\nu^{-1}\mathrm{d}k\approx\nu^{-\frac{5}{3}},$
Combining the two part and recalling the definition of $\widetilde{\mathcal{D}}$, then (3.3) follows. The proof of (3.4) is similar to (3.3). Indeed
$\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k$ $\lesssim\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\|\partial_{y}\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k$
$\lesssim\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{2}^{\frac{1}{2}}\|\omega_{k}\|_{2}^{\frac{1}{2}}\mathrm{d}k$
$\lesssim\left(\int_{\mathbb{R}}\langle k\rangle^{-2m}\gamma_{k}^{-\frac{1}{2}}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle k\rangle^{2m}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{4}}\left(\int_{\mathbb{R}}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{4}}.$
We still split the integral domain into $|k|\geq\nu^{-\frac{1}{3}}$ and $|k|<\nu^{-\frac{1}{3}}$, and recall $m>\frac{3}{4}$ and $\nu<1$, then we have
@@
$\int_{|k|<\nu^{-\frac{1}{3}}}\langle k\rangle^{-2m}\gamma_{k}^{-\frac{1}{2}}\mathrm{d}k\lesssim\int_{|k|<\nu^{-\frac{1}{3}}}\left(\nu^{-\frac{2}{3}}\right)^{-\frac{1}{2}}\mathrm{d}k\approx 1.$
Hence (3.4) follows. ∎